the area of a rectangle is 27yd^2,and the length of the rectangle is 3yd less than twice the width.Find the dimensions
Found 2 solutions by josgarithmetic, ewatrrr:
Answer by josgarithmetic(39616)

L=-3+2w and wL=27.
Substitution gives a quadratic equation, and one value will make sense but the other value will not...

Answer by ewatrrr(24785)

 
Hi,
.
A = Lw = (2w-3)w = 27yd^2
2w^2 - 3w - 27 = 0
(2w -9)(w + 3) = 0, w = 9/2yd and L = 6yd